# Quantum annealing versus classical solvers: Applications, challenges and limitations for optimisation problems

### v2

@@ 5 Conclusions
We provided an introduction to quantum annealing and conducted a benchmark of D-Wave’s current generation of quantum annealing-based hybrid solvers against state-of-the-art classical algorithms. Through a comparison of computational times and the corresponding best solutions across several case studies spanning a wide range of optimisation problems, we conclude that a state-of-the-art hybrid solver running on the most advanced quantum annealer has reached a point where it is competitive with the best classical approaches for a limited set of real-world problems.
@@
Although the hybrid solver approach enables larger problem sizes than what is currently possible for full quantum approaches, such problems still have to be heavily partitioned into many subproblems and stuck back together, leading to increasing computational time [34].
@@
By increasing the complexity of the problem through systematically adding additional constraints, we conclude that the solution quality diverges more and more from the classical solvers for an increasing number of constraints at D-Wave’s scalability limit. Additionally, the increase in run time is so drastic that the hybrid solver shows no benefit at this scale. Although we did see some competitive computational times by D-Wave when handling quadratic constraints in Section 3.3, the hybrid solver was not able to find the optimal solution with a factor of approximately two off in most runs. It is however for the BQP that D-Wave’s hybrid solver seems the most promising. Due to the mathematical structure of BQPs being so closely related to the Ising Hamiltonian of Equation 5, D-Wave’s hybrid solver performs better than all three classical solvers. The most competitive classical solver is IPOPT, but D-Wave does find slightly better solutions in a shorter amount of time. Compared with CPLEX, which due to its deterministic nature will find the optimal solution, D-Wave is far superior in regards to computational time. Regarding the MILP unit commitment problem, the benchmark underscored the importance of fine-tuning the minimum run time to enhance the solution quality. Despite various adjustments and reduced problem cases, D-Wave’s LeapHybridCQMSolver was unable to surpass Gurobi in terms of computational time or solution quality. As this specific problem has no quadratic binary instance in either the objective function or in any of the constraints, the case study indicates that these instances are currently necessary to potentially show a computational benefit at the moment. While it is promising that D-Wave can directly solve full MILP problems using their hybrid solvers, the computational advantage appears to be limited to BQP problems.
@@
The findings presented in this paper indicate that D-Wave’s hybrid solvers and Advantage quantum computers currently exhibit competitiveness with classical optimization algorithms, but this holds true for a limited range of problems. While D-Wave’s hybrid solver performs well with binary quadratic problems, its competitiveness diminishes when dealing with non-binary, non-quadratic problems, or combinations thereof. Consequently, quantum annealers with broader capabilities have strong prospects for the future. The field of quantum annealing and quantum computing, in general, is undergoing rapid and promising advancements. For instance, between 2020 to 2021 alone, D-Wave enhanced the performance of the Advantage machine, achieving an approximately eightfold increase in its ability to find solutions for Satisfiability problems and obtaining solutions for 3D lattice problems twice as fast [59]. They have already announced Zephyr for 2024 [48]. As also classical solvers are still experiencing regular performance updates, regular benchmarking research, as demonstrated in this paper, becomes imperative.
@@
The $rts\_gmlc$ dataset used for the unit commitment problem can be found in [58]. All other datasets used/and or analyzed during the current study are available from the corresponding author upon reasonable request.

### v3

@@ 5 Conclusions
We provided an introduction to quantum annealing and conducted a benchmark of D-Wave’s current generation of quantum-annealing-based hybrid solvers against state-of-the-art classical algorithms. Through a comparison of computational times and the corresponding best solutions across several case studies spanning a wide range of optimisation problems, we conclude that a state-of-the-art hybrid solver running on the most advanced quantum annealer has reached a point where it is competitive with the best classical approaches for a limited set of real-world problems.
@@
Although the hybrid-solver approach enables larger problem sizes than what is currently possible for full quantum approaches, such problems still have to be heavily partitioned into many subproblems and stuck back together, leading to increasing computational time [35].
@@
By increasing the complexity of the problem through systematically adding additional constraints, we conclude that the solution quality diverges more and more from the classical solvers for an increasing number of constraints at D-Wave’s scalability limit. Additionally, the increase in run time is so drastic that the hybrid solver shows no benefit at this scale. Although we did see some competitive computational times by D-Wave when handling quadratic constraints in Section 3.3, the hybrid solver was not able to find the optimal solution with a factor of approximately two off in most runs. It is for the BQP that D-Wave’s hybrid solver seems the most promising. Due to the mathematical structure of BQPs being so closely related to the Ising Hamiltonian of Equation 5, D-Wave’s hybrid solver performs better than all three classical solvers. The most competitive classical solver is IPOPT, but D-Wave does find slightly better solutions in a shorter amount of time. Compared with CPLEX, which due to its deterministic nature will find the optimal solution, D-Wave is superior in regards to computational time. Regarding the MILP unit commitment problem, the benchmark underscored the importance of fine-tuning the minimum run time to enhance the solution quality. Despite various adjustments and reduced problem cases, D-Wave’s LeapHybridCQMSolver was unable to surpass Gurobi in terms of computational time or solution quality. As this specific problem has no quadratic binary instance in either the objective function or in any of the constraints, the case study indicates that these instances are currently necessary to potentially show a computational benefit. While it is promising that D-Wave can directly solve full MILP problems using their hybrid solvers, the computational advantage appears to be limited to BQP problems.
@@
The findings presented in this paper indicate that D-Wave’s hybrid solvers and Advantage quantum computers currently exhibit competitiveness with classical optimisation algorithms, but this holds true only for a limited range of problems. While D-Wave’s hybrid solver performs well with binary quadratic problems, its competitiveness diminishes when dealing with non-binary, non-quadratic problems, or combinations thereof. Consequently, quantum annealers with broader capabilities have strong prospects for the future. The field of quantum annealing and quantum computing, in general, is undergoing rapid and promising advancements. Between 2020 to 2021 alone, D-Wave enhanced the performance of the Advantage machine, achieving an approximately eightfold increase in its ability to find solutions for Satisfiability problems and obtaining solutions for 3D lattice problems twice as fast [66]. They have already announced Zephyr for 2024 [54]. As classical solvers are experiencing regular performance updates, benchmarking research, as demonstrated in this paper, becomes imperative.
Acknowledgements
We acknowledge the support of the iDesignRES (Integrated Design of the Components of the Energy System to Plan the Uptake of Renewable Energy Sources)1 project which has received funding from the Horizon Europe research and innovation program under Grant Agreement No. 101095849.
@@ Acknowledgements
The $rts\_gmlc$ dataset used for the unit commitment problem can be found in [64]. All other datasets used and/or analyzed during the current study are available from the corresponding author upon reasonable request.